OpenStudy (anonymous):

is 4(-2 - -4)^4 = 64? Right?

OpenStudy (texaschic101):

4(-2 - (-4)^4 = (do parenthesis first) 4(-2 + 4)^4 = 4(2)^4 = (now exponents) 4(16) = 64 I think your right.

OpenStudy (anonymous):

:) Thanks